OpenStudy (djbreezy):

Complete the equations of the system in slope-intercept form. Use a decimal for the slope if necessary. Line 1 x y 0 3 1 5 Line 2 x y –1 1 –2 –1 Line 1: y = x + Line 2: y = x + @perl .......i have no idea

OpenStudy (djbreezy):

@perl i have no ideo

OpenStudy (anonymous):

First start with the slope formula of m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

And then for line 1 you already know the y intercept so y=mx+b should be easier to figure out

OpenStudy (djbreezy):

i havt a hard time with this

OpenStudy (anonymous):

Ok for slope formula. y2-y1=5-3 and the. You divide that by x2-x1 which equals 1-0. So the slope for line 1 is m=2/1=2.

OpenStudy (perl):

line 1 slope: $$\Large {m=\frac{5-3}{1-0}} $$

OpenStudy (djbreezy):

m=2/1?

OpenStudy (anonymous):

Once you find out the slope like perl gave a good illustration for, look at the ordered pairs. Do you have an x that is 0? Well we know that for lines the y intercept is (0,y)

OpenStudy (perl):

the y intercept is the point (0,b)

OpenStudy (perl):

thats given in the directions

OpenStudy (perl):

m = 2 y intercept b is 3 y = mx + b now put it together

OpenStudy (djbreezy):

can u give me the answer i'm completely lost now

OpenStudy (anonymous):

Wow, you have all type of K12 questions posted here. Did you know that is copyright infringement? I bet your teachers would love to know I found this....